OpenStudy (denise):

Find the slope. (-2,7) and (-1,2)

OpenStudy (sandra):

hey so do you know the slope formula?

OpenStudy (denise):

no

OpenStudy (sandra):

ok, well, to find the slope between two coordinates, you use this formula for two coordinates (x1,y1) and (x2, y2): 1. slope = (y2 - y1)/(x2 - x1)

OpenStudy (sandra):

so in the case of your problem: x1 = -2, x2 = -1, y1 = 7, and y2 = 2

OpenStudy (sandra):

so stick those into the slope formula and you will have your answer

OpenStudy (sandra):

you want to try it?

OpenStudy (denise):

It would be (-2-7) (-1-2) right?

OpenStudy (denise):

(5), (1)???

OpenStudy (sandra):

well, the slope is one number, that basically represents "for every x coordinate I move to the right, how many y coordinates do I move up?"

OpenStudy (sandra):

so the slope = (y2 - y1)/(x2 - x1) = (2 - 7)/ (-1 - (-2))

OpenStudy (sandra):

oops sorry messed up

OpenStudy (sandra):

or sorry did not mess up, the answer is -5/1 = -5

OpenStudy (denise):

through (2,1) slope =-3/2 How do I write the slope intercept form of the equation?